OpenStudy (anonymous):

A gas sample is cooled from -20.0°C to -57.0°C while the volume is increased from 2.00 L to 4.50 L. If the initial pressure is 0.122 atm, what is the final pressure?

OpenStudy (abmon98):

use combined gas law P1*V1/T1=P2*V2/T2 0.122*2.00/(273+-20)=P2*4.50/(-57+273)

OpenStudy (abmon98):

We are changing the temperature from degree celsius to kelvin by adding to 273.